: Qd = 20,000 - 3P Qs = 15,000 + 2P




Where ‘P’ is the price in rupees of a renewable energy resources, ‘Qd’ is quantity demanded for renewable energy resources, a ‘Qs’ is quantity supplied for renewable energy resources.

a) Calculate the equilibrium price and equilibrium quantity of renewable energy resources.

b) Calculate price elasticity of supply using point elasticity method when renewable energy sector is in equilibrium. Also, interpret the result

c) What will happen to the equilibrium quantity and equilibrium price of renewable energy resources if energy sector improves the technology? (Graph is not required)


1
Expert's answer
2021-12-07T08:53:43-0500

(a)"Q_ d=Q_s"

"20000-3p=15000+2p"


"5000=5p"

p = 1000

Equilibrium price=1000


Equilibrium quantity="15000+(2\times1000)"


"Q=17000"


(b)

"PES=\frac{(Qs2 – Qs1)}{[(Qs2+Qs1)/2]}"

Replacing the values in equilibrium


"PES=\frac{(17000 – 17000)}{[(17000+17000)/2]}"

PES = 0:

Supply is perfectly inelastic. There is no change in quantity if prices change. PES = infinity: Supply is perfectly elastic. An decrease in prices will lead to zero units produced.


(c)Improvement in technology in renewable energy sources will reduce the cost of renewable energy sources production shifting the supply curve downwards to the right. The demand curve will not shift but the market will move along the demand curve to a new equilibrium. Equilibrium price will fall and equilibrium quantity for renewable energy sources will increase.